OpenStudy (anonymous):

Simplify the radical expression: ^3√8 Please explain how you got this! Thanks!

OpenStudy (anonymous):

2 2 * 2 * 2 = 8

OpenStudy (anonymous):

where did you get a 2?

OpenStudy (anonymous):

The cube root of a number is the number that when cubed equals the original number, for 8 that is 2.

OpenStudy (anonymous):

i did not get any of that!

OpenStudy (anonymous):

Wait, did you mean \[\sqrt[3]{8}\] OR \[\sqrt{8}\]?

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

Okay, that is 2.

OpenStudy (anonymous):

The Cube Root of 8 is 2.

OpenStudy (anonymous):

ok I see

OpenStudy (anonymous):

so the answer is just 2?

OpenStudy (anonymous):

Yes, just "2"

OpenStudy (phi):

BTW, there is no trivial way to find the cube root. You typically solve these by memorizing the first few cubes. 1 1 2 8 3 27 4 64 5 125 10 1000 or use a calculator. Later on you might learn about logarithms.